Someone Help Me Please? :)
Vito invested $500, part at 9% per annum, and the rest at 11% per annum. After one year, the interest earned on the 9% investment was $20 less than the interest on the 11% investment. How much did he invest at each rate?

Answer by stanbon(75887)

Vito invested $500, part at 9% per annum, and the rest at 11% per annum. After one year, the interest earned on the 9% investment was $20 less than the interest on the 11% investment. How much did he invest at each rate?
--------------------
Let amount invested at 9% be "x" ; Interest on this is 0.09x dollars
----------------
Amount invested at 11% is "500-x" ; Interest on this is 0.11(500-x)=55-0.11x
------------------
EQUATION:
9% int. = 11% - 20
0.09x = 55-0.11x -20
0.20x = 35
x = $175.00 (Amt invested at 9%)
500-x = $325 (Amt invested at 11%)
=================
Cheers,
Stan H.